The gravitational force exerted on a baseball is 2.15 N down. A pitcher throws the ball horizontally with velocity 12.5 m/s by uniformly accelerating it along a straight horizontal line for a time interval of 190 ms. The ball starts from rest.
1
Expert's answer
2020-06-25T09:35:17-0400
"d=0.5vt=0.5(12.5)(0.19)=1.19\ m"

"F=ma=m\frac{v}{t}=\frac{Wv}{gt}"

"F=\frac{(2.15)(12.5)}{(9.8)(0.19)}=14.4\ N"